### v1

@@ Abstract
Latent diffusion models have recently demonstrated superior capabilities in many downstream image synthesis tasks. However, customization of latent diffusion models using unauthorized data can severely compromise the privacy and intellectual property rights of data owners. Adversarial examples as protective perturbations have been developed to defend against unauthorized data usage by introducing imperceptible noise to customization samples, preventing diffusion models from effectively learning them. In this paper, we first reveal that the primary reason adversarial examples are effective as protective perturbations in latent diffusion models is the distortion of their latent representations, as demonstrated through qualitative and quantitative experiments. We then propose the Contrastive Adversarial Training (CAT) utilizing adapters as an adaptive attack against these protection methods, highlighting their lack of robustness. Extensive experiments demonstrate that our CAT method significantly reduces the effectiveness of protective perturbations in customization configurations, urging the community to reconsider and enhance the robustness of existing protective perturbation methods. Code is available at https://github.com/senp98/CAT.
@@ 1 Introduction
Diffusion-based image generation models, especially the Latent Diffusion Models (LDMs) Rombach et al. (2022), have exhibited strong capabilities across various downstream image generation tasks, including text-to-image synthesis Dhariwal & Nichol (2021); Rombach et al. (2022), image-to-image translation Saharia et al. (2022); Zhang et al. (2023), image editing Kawar et al. (2023), and image inpainting Lugmayr et al. (2022). By utilizing the open-source LDM such as Stable Diffusion Stability.ai (2024) as the pre-trained model, users can customize image generation models on their datasets at a significantly lower cost than training from scratch. Despite their prominence and popularity, the customization of LDMs also raises significant concerns regarding privacy and Intellectual Property Rights (IPRs) Dixit (2023); LLP (2023). It can be misused to generate fake human face images of a specific identity Li et al. (2022) or mimic the style of a specific artist by utilizing unauthorized data. Undoubtedly, such malicious customization severely compromises the rights of data owners and raises ethical concerns across the community.
@@
Several protective perturbation methods Salman et al. (2023); Liang et al. (2023); Shan et al. (2023); Van Le et al. (2023); Liang & Wu (2023); Xue et al. (2023); Liu et al. (2024) have been proposed to address these concerns in diffusion-based image generation. These methods introduce imperceptible perturbations to image samples in the customization dataset, creating adversarial examples that effectively degrade the performance of downstream task customization. Although significant progress has been made in developing perturbation methods, the robustness of these protection techniques has received limited attention. Recent studies Zhao et al. (2024); Cao et al. (2024); Hönig et al. (2024) have investigated the resilience of protective perturbations against purification-based adaptive attacks. By purifying the protected images before customization, these methods effectively compromise the robustness of existing protective perturbation. However, existing purification-based adaptive attacks depend on optimization, resulting in high computational costs. Without an efficient mechanism to detect the protected perturbed images, purification must be applied to every sample in the customization dataset, increasing resource demand. Furthermore, purification without prior knowledge of the original data distribution introduces uncertainty, potentially altering the learned distribution. In extreme cases, optimization-based purification may entirely replace the original image content if the substituted one achieves a lower loss for the optimization objective.
@@
In this paper, we introduce Contrastive Adversarial Training (CAT) from a novel perspective, focusing on model adaptation rather than purification as a solution approach. We first reveal that the primary reason adversarial examples are effective as protective perturbations in LDM customization is that their latent representations are distorted. Building on this insight, we develop the CAT method targeting the latent autoencoder of the LDM, utilizing the contrastive adversarial loss to minimize the reconstruction loss of protected image samples. During adversarial training, we introduce LoRA adapters Hu et al. (2022) to specific layers of the latent autoencoder, ensuring that only the adapter weights are modified. This approach expands the parameter space to mitigate the effectiveness of adversarial examples and reduces computational costs, enhancing its adaptability across different scenarios. Extensive experiments demonstrate that applying CAT decreases the effectiveness of protective perturbations, urging the community to reconsider their robustness. In summary, our main contributions are as follows:
@@ 2.1 Diffusion-based Text-to-Image Models
Diffusion-based text-to-image models have demonstrated remarkable capabilities across various downstream tasks. Images are learned through the forward diffusion process, while generation involves denoising in the reverse diffusion process, typically using a U-Net Ronneberger et al. (2015) as the backbone model Ho et al. (2020). Latent diffusion models Rombach et al. (2022) improve training and inference efficiency while maintaining generation quality by encoding samples into their latent space representations through a latent autoencoder before the diffusion process. Stable Diffusion Stability.ai (2024) is later released as an open-source LDM, significantly boosting the development of various downstream applications in diffusion-based text-to-image generation.
@@ 2.2 Customization of LDMs
Customization of LDMs can be applied to various downstream tasks, which are classified into two categories: text-driven image synthesis and text-driven image editing. Text-driven image synthesis can be divided into (1) object-driven image synthesis and (2) style mimicry. The former generates images of a specific object based on the given text prompt, while the latter generates images that adopt a specific style defined by the text prompt. DreamBooth Ruiz et al. (2023) customizes text-to-image models by fine-tuning the U-Net and text encoder with a few images and employing a class-specific prior preservation loss, enabling the image generation of the target subject while preserving its key features. LoRA Hu et al. (2022) was initially proposed to reduce the cost of fine-tuning large language models and later applied to diffusion model customization. It assumes that the model’s weight updates form a low-rank matrix, which can be represented through its rank decomposition. By introducing lightweight adaptation layers to a frozen pre-trained model and updating only these layers during customization, LoRA significantly lowers the computational cost during customization.
@@ 2.3 Protective Perturbations
The idea of protective perturbations originates from adversarial attacks Goodfellow et al. (2014) in classification models, where small perturbations mislead the model into making incorrect predictions. Recent work Salman et al. (2023); Liang et al. (2023); Shan et al. (2023); Van Le et al. (2023); Liang & Wu (2023); Xue et al. (2023) shows that diffusion-based generative models are also vulnerable to such attacks, with adversarial examples serving as protective perturbations that significantly degrade the performance of downstream tasks, preventing unauthorized customization. Photoguard Salman et al. (2023) first introduced protective perturbations to increase the cost of diffusion-based image editing such as SDEdit Meng et al. (2021). The perturbation is created by maximizing the latent and diffusion denoising loss over the protected images. Both AdvDM Liang et al. (2023) and SDS Xue et al. (2023) generate perturbations targeting the entire LDM through end-to-end optimization to prevent learning, while Glaze Shan et al. (2023) focuses on perturbations against the latent encoder of the LDM. Mist Liang & Wu (2023) combines latent and denoising losses in perturbation generation, while Anti-DreamBooth Van Le et al. (2023) creates protective perturbations against the DreamBooth Ruiz et al. (2023) customization approach. MetaCloak Liu et al. (2024) leverages a meta-learning framework with transformation sampling to generate transferable and robust perturbations.
Purification-based adaptive attacks have recently been proposed against existing protection methods. IMPRESS Cao et al. (2024) introduces a purification technique that minimizes the difference between the protected image and its reconstruction from the latent autoencoder of the LDM. GrIDPure Zhao et al. (2024) builds upon DiffPure Nie et al. (2022) to pre-process protected samples. Additionally, the authors in Hönig et al. (2024) propose using Noisy-Upscaling to purify protected image samples before customization. However, all these methods decrease the effectiveness of protective perturbations by purifying the protected images. To the best of our knowledge, we are the first to address the robustness of protection perturbations in the customization of LDMs from the novel perspective of model adaptation.
@@ 3.1 Distortion in Latent Representations
We first identify a key reason why adversarial examples serve as effective protective perturbations in LDMs: their representations are distorted in the latent space. As illustrated in Figure 2, the latent autoencoder of a LDM first encodes a given clean image $x_{c}$ from pixel space into a latent representation $z_{c}$. Subsequently, the denoising module takes the latent representation as input because it has significantly lower dimensionality while still preserving the semantic features. Additionally, it requires that the autoencoder can correctly decode the latent representation $z_{c}$ back to pixel space, ensuring that the reconstructed image $\tilde{x}_{c}$ matches the original image $x_{c}$ in the ideal case. However, the mappings between pixel and latent space are distorted for adversarial examples. For an adversarial example $x_{a}$ within a perturbation budget $\delta$, the latent representation $\hat{z}_{a}$ should remain close to $z_{c}$ within a distance $\delta_{z}$ and correctly decode back to $x_{a}$ in pixel space in the ideal case. However, $x_{a}$ is instead mapped to a distorted latent representation $z_{a}$. This distorted representation is decoded to $\tilde{x}_{a}$, which exhibits a significantly larger reconstruction error.
@@
Similar observations about the latent representation are also reported in previous works Xue et al. (2023); Cao et al. (2024). In contrast, our study conducts more extensive qualitative and quantitative experiments to thoroughly validate the distortion in latent representations of protected samples. To illustrate this, we use t-SNE to visualize the latent representations of various protected samples, as shown in Figure 3.
@@
Figure 3: Visualization of latent representations of clean sample $z_{c}$, noisy sample $z_{r}$, protected sample $z_{a}$, and the latent representation of protected sample $z_{a}^{(\text{cat})}$ encoded by the latent encoder with CAT adapter. Protected samples are generated using nine protective perturbation methods, with four clean images from the VGGFace2 dataset as the base. The perturbation budget for protected and noisy samples is fixed at $\delta=16/255$.
@@
We use four clean images of an identity from the VGGFace2 Cao et al. (2018) dataset as an illustration. For each clean sample $x_{c}$, we generate its adversarial example $x_{a}$ as the protected image using nine different protective perturbation methods, within the perturbation budget fixed at $\delta=16/255$. The protected sample is then encoded into the latent space through the latent encoder as $z_{a}=\mathcal{E}(x_{a})$, while the clean sample is encoded as $z_{c}=\mathcal{E}(x_{c})$. Additionally, we construct a noisy sample $x_{r}=x_{c}+clip(r,\delta)$ by adding Gaussian noise $r\sim N(0,\delta^{2})$, clipped using the same perturbation budget $\delta$. This noisy sample is encoded as $z_{r}=\mathcal{E}(x_{r})$. Finally, the adversarial sample $x_{a}$ is also encoded using the latent encoder with the CAT adapter (in the CAT-both setting), resulting in $z_{a}^{(\text{cat})}=\mathcal{E}^{(\text{cat})}(x_{a})$. For each protection method shown in the subplots, $z_{a}$, $z_{r}$ and $z_{a}^{(\text{cat})}$ are clustered with their corresponding clean sample $z_{c}$. Since four images are evaluated for each protection method, their latent representations naturally form four distinct groups. Within each group, it can be observed that the latent representation of the protected sample $z_{a}$ exhibits a larger distance from the clean sample’s latent representation $z_{c}$ compared to the noisy sample $z_{r}$ under the same perturbation budget. This pattern holds across all four image samples for every protective perturbation method, indicating that the latent representations of adversarial examples as protected samples are distorted across all evaluated protection methods.
@@
In addition to the qualitative experiments, we provide a quantitative analysis of the distances of latent representations. We define the MAE distance between the latent representation of a sample and its corresponding clean sample as $d$. Specifically, the distance between $z_{a}$ and its corresponding $z_{c}$ is denoted as $d_{a}=\|z_{a}-z_{c}\|_{1}$, while the distance between $z_{a}^{(\text{cat})}$ and its corresponding $z_{c}$ is represented as $d_{a}^{(\text{cat})}=\|z_{a}^{(\text{cat})}-z_{c}\|_{1}$. The distance between a noisy sample $z_{r}$ and its corresponding $z_{c}$ is denoted as $d_{a}=\|z_{a}-z_{c}\|_{1}$. For comparison, we present the differences between $d_{a}$ and $d_{a}^{(\text{cat})}$ over $d_{r}$, as shown in Figure 4. As we can observe, $d_{a}$ is larger than $d_{r}$ across all the evaluated protection methods, indicating that adversarial examples significantly distort their latent representations compared to noisy samples with the same perturbation budget. Furthermore, our proposed CAT adapter method can effectively reduce this distortion, realigning the latent representations of protected examples closer to their correct positions.

### v2

@@ Abstract
Latent diffusion models have recently demonstrated superior capabilities in many downstream image synthesis tasks. However, customization of latent diffusion models using unauthorized data can severely compromise the privacy and intellectual property rights of data owners. Adversarial examples as protective perturbations have been developed to defend against unauthorized data usage by introducing imperceptible noise to customization samples, preventing diffusion models from effectively learning them. In this paper, we first reveal that the primary reason adversarial examples are effective as protective perturbations in latent diffusion models is the distortion of their latent representations, as demonstrated through qualitative and quantitative experiments. We then propose the Contrastive Adversarial Training (CAT) utilizing lightweight adapters as an adaptive attack against these protection methods, highlighting their lack of robustness. Extensive experiments demonstrate that our CAT method significantly reduces the effectiveness of protective perturbations in customization, urging the community to reconsider and improve the robustness of existing protective perturbations. The code is available at https://github.com/senp98/CAT.
latent diffusion models, protective perturbations, adversarial training
@@ 1 Introduction
Diffusion-based image generation models, especially the Latent Diffusion Models (LDMs) (Rombach et al., 2022), have exhibited strong capabilities across various downstream image generation tasks, including text-to-image synthesis (Dhariwal & Nichol, 2021; Rombach et al., 2022), image-to-image translation (Saharia et al., 2022; Zhang et al., 2023), image editing (Kawar et al., 2023), and image inpainting (Lugmayr et al., 2022). By utilizing the open-source LDM such as Stable Diffusion (Stability.ai, 2024) as the pre-trained model, users can customize image generation models on their datasets at a significantly lower cost than training from scratch. Despite their prominence and popularity, the customization of LDMs also raises significant concerns regarding privacy and intellectual property rights (Dixit, 2023; LLP, 2023). It can be misused to generate fake human face images of a specific identity (Li et al., 2022) or mimic the style of a specific artist by utilizing unauthorized data. Undoubtedly, such malicious customization severely compromises the rights of data owners and raises ethical concerns across the community.
@@
Several protective perturbation methods (Salman et al., 2023; Liang et al., 2023; Shan et al., 2023; Le et al., 2023; Liang & Wu, 2023; Xue et al., 2024; Liu et al., 2024) have been proposed to address these concerns. These methods introduce imperceptible perturbations to image samples in the customization dataset, creating adversarial examples that effectively degrade the performance of downstream customization tasks. Although significant progress has been made in developing perturbation methods, the robustness of these protection techniques has received limited attention. Recent studies (Zhao et al., 2024; Cao et al., 2024; Hönig et al., 2025) have investigated the resilience of protective perturbations against purification-based adaptive attacks. By purifying the protected images before customization, these methods effectively compromise the robustness of existing protective perturbation. However, existing purification-based adaptive attacks depend on optimization, resulting in high computational costs. Without an efficient mechanism to detect the protected perturbed images, purification must be applied to every sample in the customization dataset, increasing resource demand. Furthermore, purification without prior knowledge of the original data distribution introduces uncertainty, potentially altering the learned distribution. In extreme cases, optimization-based purification may entirely replace the original image content if the substituted one achieves a lower loss for the optimization objective.
@@
In this paper, we introduce Contrastive Adversarial Training (CAT) from a novel perspective, focusing on model adaptation rather than purification as an adaptive attack. We first reveal that the primary reason adversarial examples are effective as protective perturbations in LDM customization is that their latent representations are distorted. Building on this insight, we develop the CAT method targeting the latent autoencoder of the LDM, utilizing the contrastive adversarial loss to minimize the reconstruction loss of protected image samples. During adversarial training, we introduce LoRA adapters (Hu et al., 2022) to specific layers of the latent autoencoder, ensuring that only the adapter weights are modified. This approach expands the parameter space to mitigate the effectiveness of adversarial examples and reduces computational costs, enhancing its adaptability across different scenarios. Extensive experiments demonstrate that applying CAT decreases the effectiveness of protective perturbations, urging the community to reconsider their robustness. In summary, our main contributions are as follows:
@@ 2.1 Diffusion-based Text-to-Image Models
Diffusion-based text-to-image models have demonstrated remarkable capabilities across various downstream tasks. They learn to generate images by simulating a forward diffusion process that adds noise, and a reverse diffusion process that removes it, typically using a U-Net (Ronneberger et al., 2015) as the backbone (Ho et al., 2020). Latent diffusion models (Rombach et al., 2022) improve training and inference efficiency while preserving generation quality by encoding samples into their latent representations via an autoencoder. Stable Diffusion (Stability.ai, 2024) was later released as an open-source LDM, which significantly boosted the development of various downstream applications in diffusion-based text-to-image generation.
@@ 2.2 Customization of LDMs
LDMs can be customized for a range of downstream tasks, which broadly fall into text-driven image synthesis and editing. Text-driven image synthesis is typically categorized into object-driven image synthesis and style mimicry (Peng et al., 2024). The former focuses on generating images containing specific objects described by the text prompt, while the latter aims to generate images that reflect the style specified in the text prompt. DreamBooth (Ruiz et al., 2023) customizes text-to-image diffusion models by fine-tuning both the U-Net and text encoder using a few target subject images. It introduces a class-specific prior preservation loss to prevent overfitting on the subject, enabling the model to generate diverse class images while preserving key visual features of the subject. LoRA (Hu et al., 2022), originally developed to enable parameter-efficient fine-tuning large language models, has been adapted to diffusion models to enable customization via low-rank adaptation. By introducing lightweight trainable adaptation layers into a frozen pre-trained model and updating only these layers, LoRA significantly reduces the memory and computational overhead during model customization.
@@ 2.3 Protective Perturbations
Protective perturbations build upon the idea of adversarial attacks (Goodfellow et al., 2015) in classification models, which introduce imperceptible input perturbations to mislead models into incorrect predictions. Recent studies (Salman et al., 2023; Liang et al., 2023; Shan et al., 2023; Le et al., 2023; Liang & Wu, 2023; Xue et al., 2024) have shown that diffusion-based generative models are also vulnerable to such attacks. In this context, adversarial examples serve as protective perturbations that significantly degrade the performance of downstream tasks and prevent unauthorized customization. Photoguard (Salman et al., 2023) first introduces protective perturbations to increase the cost of diffusion-based image editing methods such as SDEdit (Meng et al., 2022). The perturbation is generated by jointly maximizing the latent reconstruction loss and the diffusion denoising loss with respect to the protected images. AdvDM (Liang et al., 2023) and SDS (Xue et al., 2024) generate perturbations that target the entire LDM through end-to-end optimization to prevent unauthorized model customization. In contrast, Glaze (Shan et al., 2023) focuses on attacking the latent encoder of LDMs to distort the encoded representations. Mist (Liang & Wu, 2023) optimizes the joint latent and denoising losses to generate perturbations, while Anti-DreamBooth (Le et al., 2023) specifically targets the DreamBooth (Ruiz et al., 2023) customization approach. MetaCloak (Liu et al., 2024) leverages a meta-learning framework with transformation sampling to generate transferable and robust perturbations.
Purification-based adaptive attacks have recently been proposed against existing protective perturbations. IMPRESS (Cao et al., 2024) introduces a purification technique that minimizes the difference between the protected image and its reconstruction from the latent autoencoder. GrIDPure (Zhao et al., 2024) extends DiffPure (Nie et al., 2022) to purify protected samples before customization, while Noisy-Upscaling (Hönig et al., 2025) achieves a similar goal using other generative models. However, all these methods decrease the effectiveness of protective perturbations through purification-based adaptive attacks. To the best of our knowledge, this is the first work to investigate the robustness of such perturbations in the context of LDM customization from the novel perspective of model adaptation. Recent RobustCLIP (Schlarmann et al., 2024) enhances the robustness of CLIP-based vision encoders through adversarial training, particularly against perturbations that disrupt text-image semantic alignment. While effective, it requires fine-tuning the entire encoder on a large set of adversarial examples to achieve general robustness. In contrast, CAT introduces lightweight adapters during customization on protected data. These adapters are detachable at inference time, leaving the original model performance unaffected.
@@ 3 Understanding Protective Perturbations
3.1 Threat Model
We first introduce the threat model considered in this work. We assume the data owner holds clean data samples and intends to share them publicly. To prevent unauthorized customization of the generative model, the owner applies a protective perturbation to each image before release and then publicly shares the perturbed dataset. In this setting, the adversary only has access to the protected data and is unaware of the specific protection method used by the data owner. We want to demonstrate that, even without knowledge of the protection technique and with access only to the protected data, the adversary can still successfully customize the LDM using our proposed CAT.
@@ 3.2 Distortion in Latent Representations
We first identify a key reason why adversarial examples serve as effective protective perturbations in LDMs: their representations are distorted in the latent space. As illustrated in Figure 2, the latent autoencoder of a LDM first encodes a given clean image $x_{c}$ from pixel space into a latent representation $z_{c}$. Subsequently, the denoising module takes the latent representation as input because it has significantly lower dimensionality while still preserving the semantic features. Additionally, it requires that the autoencoder can correctly decode the latent representation $z_{c}$ back to pixel space, ensuring that the reconstructed image $\tilde{x}_{c}$ matches the original image $x_{c}$ in the ideal case. However, the mappings between pixel and latent space are distorted for adversarial examples. For an adversarial example $x_{a}$ within a perturbation budget $\delta_{x}$, the latent representation $\hat{z}_{a}$ should remain close to $z_{c}$ within a distance $\delta_{z}$ and correctly decode back to $x_{a}$ in pixel space in the ideal case. However, $x_{a}$ is instead mapped to a distorted latent representation $z_{a}$. This distorted representation is decoded to $\tilde{x}_{a}$, which exhibits a significantly larger reconstruction error.
@@
Similar observations about the latent representation are also reported in previous works (Xue et al., 2024; Cao et al., 2024). In contrast, our study conducts more extensive qualitative and quantitative experiments to thoroughly validate the distortion in latent representations of protected samples. To illustrate this, we visualize the latent representations of various protected samples using t-SNE (der Maaten & Hinton, 2008), as shown in Figure 3. We also include the UMAP visualization (McInnes et al., 2018) in Figure 8, which is provided in Appendix A, to enhance the relative distance demonstration.
@@
Figure 3: Visualization of latent representations using t-SNE of clean sample $z_{c}$, noisy sample $z_{r}$, protected sample $z_{a}$, and the latent representation of protected sample $z_{a}^{(\text{cat})}$ encoded by the latent encoder with CAT adapter. Protected samples are generated using nine protective perturbation methods, with four clean images from the VGGFace2 dataset as the base. The perturbation budget for protected and noisy samples is fixed at $\delta_{x}=16/255$. Each number labels the representations obtained from the same image.
@@
We use four clean images of an identity from the VGGFace2 (Cao et al., 2018) dataset as an illustration. For each clean sample $x_{c}$, we generate its adversarial example $x_{a}$ as the protected image using nine different protective perturbation methods, within the perturbation budget fixed at $\delta_{x}=16/255$. The protected sample is then encoded into the latent space through the encoder as $z_{a}=\mathcal{E}(x_{a})$, while the clean sample is encoded as $z_{c}=\mathcal{E}(x_{c})$. Additionally, we construct a noisy sample $x_{r}=x_{c}+clip(r,\delta_{x})$ by adding Gaussian noise $r\sim N(0,\delta_{x}^{2})$, clipped using the same perturbation budget $\delta_{x}$. This noisy sample is encoded as $z_{r}=\mathcal{E}(x_{r})$. Finally, the adversarial sample $x_{a}$ is also encoded using the latent encoder with the CAT adapter (in the CAT-both setting), resulting in $z_{a}^{(\text{cat})}=\mathcal{E}^{(\text{cat})}(x_{a})$. For each protection method shown in the subplots, $z_{a}$, $z_{r}$ and $z_{a}^{(\text{cat})}$ are clustered with their corresponding clean sample $z_{c}$. Since four images are evaluated for each protection method, their latent representations naturally form four distinct groups. Within each group, it can be observed that the latent representation of the protected sample $z_{a}$ exhibits a larger distance from the clean sample’s latent representation $z_{c}$ compared to the noisy sample $z_{r}$ under the same perturbation budget. This pattern holds across all four image samples for every protective perturbation method, indicating that the latent representations of adversarial examples as protected samples are distorted across all evaluated protection methods. This observation is also supported by the UMAP visualization.
@@
In addition to qualitative experiments, we conduct a quantitative analysis of latent representation distances to support our findings. We define the MAE distance between the latent representation of a sample and its corresponding clean sample as $d$. Specifically, the distance between $z_{a}$ and its corresponding $z_{c}$ is denoted as $d_{a}=\|z_{a}-z_{c}\|_{1}$, while the distance between $z_{a}^{(\text{cat})}$ and its corresponding $z_{c}$ is represented as $d_{a}^{(\text{cat})}=\|z_{a}^{(\text{cat})}-z_{c}\|_{1}$. The distance between a noisy sample $z_{r}$ and its corresponding $z_{c}$ is denoted as $d_{r}=\|z_{r}-z_{c}\|_{1}$. For comparison, we present the relative differences between $d_{a}$ and $d_{a}^{(\text{cat})}$ over $d_{r}$, as shown in Figure 4. We observe that $d_{a}$ is consistently larger than $d_{r}$ across all evaluated protection methods, indicating that adversarial examples significantly distort their latent representations compared to noisy samples with the same perturbation budget. Furthermore, our proposed CAT adapter effectively reduces this distortion, realigning the latent representations of protected examples toward their original positions.